# Simultaneous Inference in Multiple Matrix-Variate Graphs for High-Dimensional Neural Recordings

### v1

@@ 4 Theoretical Properties
We make the following assumptions on the observed dataset.
@@
For $l=1,\dots,m$, let $\{\lambda^{(\mathcal{T},l)}_{i}\}_{i=1,\dots,p}$ are the eigenvalues of $\Sigma^{(\mathcal{T},l)}$ while $\frac{1}{\kappa_{3}}\leq\lambda^{(\mathcal{T},l)}_{1}\leq\lambda^{(\mathcal{T},l)}_{2}\leq\dots\leq\lambda^{(\mathcal{T},l)}_{p}\leq\kappa_{3}$ for some constant $\kappa_{3}>0$; define and assume $\{\lambda^{(\mathcal{S},l)}_{i}\}_{i=1,\dots,q}$ similarly for $\Sigma^{(\mathcal{S},l)}$.
@@
Let $d$ be the group-wise maximum node degree, i.e.,
$d:=\max_{i}\left\lvert\left\{j\in[p]\backslash\{i\}:\Omega^{(\mathcal{S},l)}_{ij}\neq 0\text{ for some }l\in[m]\right\}\right\rvert.$
We assume group sparsity of the partial-correlation graph in spatial association by $d\cdot\frac{\max\{m,\log(mn_{0}pq)\}}{(mn_{0}p)^{1/2}}\rightarrow 0$ as $n_{0}\rightarrow\infty$.
@@
We assume the temporal precision matrix $\Omega^{(\mathcal{T},l)}$ for each $l=1,\dots,m$ has Cholesky decomposition as in Eq. 9 where $\beta^{(\mathcal{T},l)}$ satisfies $\lvert\beta^{(\mathcal{T},l)}_{st}\rvert<\kappa_{5}(t-s)^{-\alpha_{l}-1}$ for any $t$ and $s$ such that $s<t$ and some $\alpha_{l}>0$. We further assume that, for $\alpha_{0}=\min_{l=1,\dots,m}\alpha_{l}$, $\frac{\log(mn_{0}pq)}{(n_{0}q)^{1-1/(\alpha_{0}+1)}}\rightarrow 0$ as $n_{0}\rightarrow\infty$.
Assumption 1 assumes that the sample sizes across graphs are balanced, with $n_{0}$ representing this common level. Assumption 2 ensures identifiability. Assumption 3 is a standard assumption on eigenvalues commonly used in covariance estimation and graphical models (Cai et al., 2016b, ). Assumption 4 assumes that the spatial (column) precision matrices are sparse, and it imposes a constraint on the spatial dimension relative to the number of samples, the temporal dimension, and the number of graphs. The first part of Assumption 5 is reasonable for neural time series, as neural data, particularly LFPs, are often modeled as a low-order autoregressive process, a common assumption in the literature (Bickel and Levina,, 2008; Liu and Ren,, 2020). The second part, similar to Assumption 4, restricts the temporal dimension.
@@ 4.1 Non-asymptotic error bound for the group Lasso estimate
We first provide a theoretical justification for our group Lasso procedure proposed in Section 3.1. Although with correlated rows, our results below demonstrate that the optimal convergence rates for the estimation error, defined as $\Delta^{(\mathcal{S},l)}_{\cdot,i}:=\widehat{\beta}^{(\mathcal{S},l)}_{\cdot,i}-{\beta}^{(\mathcal{S},l)}_{\cdot,i}$, and prediction can be still obtained compared to the case with i.i.d. samples. The proof is provided in Appendix B.2.
@@ 4.2 Theoretical justification for the simultaneous edge testing
We now present the theoretical results for the simultaneous test in Section 3.2. In the proof of Proposition 4.2, we show that the error in the partial correlation estimate $\widehat{\rho}^{(\mathcal{S},l)}_{ij}$ from Eq. 4 is dominated by the leading term
@@
where $\widetilde{\phi}_{ij}^{(\mathcal{S},l)}:=\widetilde{\Phi}^{(\mathcal{S},l)}_{ij}-\Phi^{(\mathcal{S},l)}_{ij}$ is the empirical error, given by $\widetilde{\phi}_{ij}^{(\mathcal{S},l)}=\frac{1}{n_{l}p}\sum_{k=1}^{n_{l}}\epsilon^{(\mathcal{S},k,l)\top}_{\cdot,i}\epsilon^{(\mathcal{S},k,l)}_{\cdot,j}-\Phi^{(\mathcal{S},l)}_{ij}.$ Since the leading error term $\Theta_{ij}^{(\mathcal{S},l)}$ is a linear functional of $\widetilde{\phi}^{(\mathcal{S},l)}_{ij}$, which is an average over $t=1,\dots,p$ and $k=1,\dots,n_{l}$, the central limit theorem ensures that the partial correlation estimate $\widehat{\rho}^{(\mathcal{S},l)}_{ij}$ and the single-edge test statistic $\widehat{T}_{ij}$ converge in distribution to Gaussian limits for each edge $(i,j)$.
For the multiple-edge test statistic $\lVert\widehat{T}_{E}\rVert_{\infty}$, the Gaussian approximation error is related to Berry–Esseen bounds over hyper-rectangles. The seminal work of Chernozhukov et al., (2013) established such bounds under high-dimensional settings, where the dimension exceeds the sample size. Subsequent works, as reviewed by Chernozhukov et al., 2023a , have focused on improving the convergence rates. Recently, Chernozhukov et al., 2023b achieved a near-optimal rate, which we use to derive a sharp Gaussian approximation error bound for $\lVert\widehat{T}_{E}\rVert_{\infty}$. The proof is detailed in Appendix B.3.
@@
The following theorem mirrors the previous proposition, with the key difference being that we replace the population covariance $S_{EE}$ with its plug-in estimator $\widehat{S}_{EE}$. In essence, as long as $S_{EE}$ is well-estimated under the $\lVert\cdot\rVert_{\infty}$-norm, the Gaussian approximation results remain valid. Using the convergence rate of $\widehat{S}_{EE}$ in the $\lVert\cdot\rVert_{\infty}$-norm, we derive the following bootstrap theorem, based on the relationship between convergence rate and bootstrap error as given in Lemma 2.1 of Chernozhukov et al., 2023b .
@@
$\sup_{x>0}\left\lvert\mathbb{P}[\lVert\widehat{T}_{E}-T_{E}\rVert_{\infty}>x]-\mathbb{P}[\lVert\widehat{Z}\rVert_{\infty}>x|\mathcal{D}]\right\rvert$
@@
The above theorem establishes the theoretical foundation for the simultaneous multiple edge testing procedure in Algorithm 1. Next, we formally state the validity of our testing procedure as well as a power analysis. The proof is given in Appendix B.5.
@@
[…]
Remark 4.5.
The theorem implies that as long as the sum of the partial correlations $\rho^{(\mathcal{S},l)}_{ij}$ across $l$, i.e., $\sum_{l=1}^{m}\lvert\rho^{(\mathcal{S},l)}_{ij}\rvert$, is above the order of $\sqrt{m\log q/(n_{0}p)}$ for some $(i,j)\in E$, the power converges to $1$ as $n_{0}\rightarrow\infty$. Assuming the same order of $\rho^{(\mathcal{S},l)}_{ij}$ across sessions, the power of the test converges to $1$ as $n_{0}\rightarrow\infty$, if $\max_{(i,j)\in E}\lvert\rho^{(\mathcal{S},l)}_{ij}\rvert$ is larger than $\sqrt{\log q/(mn_{0}p)}$. In contrast, the corresponding detection boundary is $\sqrt{\log q/(n_{0}p)}$ if we do not aggregate multiple graphs. Therefore, by borrowing the information from multiple graphs or sessions, we can reduce the detection accuracy by a factor of root $m$. In addition, when the temporal covariance matrix $\Sigma^{(\mathcal{T},l)}=I$ for all $l=1,...,m$, the model reduces to the multiple vector-variate Gaussian graphical model, as studied in Ren et al., (2019), with an effective sample size $n_{0}p$. According to their Theorem 2.3, the optimal separating rate for detecting a single edge, in terms of the sum of the partial correlations, is $\sqrt{m/(n_{0}p)}$. This suggests that our linear functional-based test is nearly optimal in achieving the testable region boundary.
4.3 Non-asymptotic error bound for the temporal covariance matrix estimate
In this section, we present the estimation error bounds for the session-specific temporal covariance and precision matrices obtained in Section 3.3. While results for i.i.d. samples are available in Liu and Ren, (2020), no such results exist for the correlated samples considered in our model. Therefore, we provide a self-contained analysis, which may be of independent interest. The proof is detailed in Appendix B.7.
Proposition 4.6.
Theorem 4.7.
Consequently, the Frobenius norms of $\Sigma^{(\mathcal{T},l)}$ can be consistently estimated, which is sufficient for our main result Theorem 4.3. The proof is given in Appendix B.8.
Corollary 4.8.
for sufficiently large $n_{0}$.
@@ 5 Numerical Studies
In Section 5.1, we present simulation studies to validate our theoretical results of Section 4 and evaluate the performance of our matrix-variate graphical models in comparison to several baseline approaches. In Section 5.2, we apply our method to neural recordings from two Utah arrays implanted in monkey brains during a spatial working memory task. The implementation of our proposed method is available in the mmge R package, along with vignettes for reproducing our results, accessible at github.com/HeejongBong/mmge.
@@ 5.1 Simulation Studies
We evaluated the performance of our method under three spatial graph structures, illustrated in Figure 1: (1) a random graph, where edges between nodes were generated with probability $\sqrt{\frac{3}{q}}$ for each pair $(i,j)$; (2) a hub graph, where the nodes were divided into $\lceil\frac{q}{20}\rceil$ hub groups; and (3) a chain graph. Given each spatial graph structure, we generated the spatial precision matrix by sampling the non-zero entries from $\text{Unif}(0,\frac{0.3}{2^{l-1}})$, independently across sessions $l=1,\dots,m$. For the temporal precision matrix, we generated $\Sigma^{(\mathcal{T},l)}$ using Eq. 9. Following Assumption 5, we set $\beta^{(\mathcal{T},l)}_{st}=\kappa_{5}(t-s)^{-\alpha_{l}-1}$ for $1\leq s<t\leq p$, with $\kappa_{5}=0.2$ and $\alpha_{l}=1$. For $\Phi^{(\mathcal{T},l)}$, we used the $p\times p$ identity matrix for all $l=1,\dots,m$.
In Section 5.1.1, we compare our method’s performance with existing methods in terms of edge detection and precision matrix estimation. In Section 5.1.2, we assess the coverage of the proposed bootstrap confidence region for $\lVert T_{E}\rVert_{\infty}$.
@@ 5.1.1 Edge-wise Estimation Comparison
  • (M1): matrix-variate Gaussian multi-graph estimation method by Zhu and Li, (2018)
  • (M2): regression-based Gaussian graphical model estimation method by Ren et al., (2019)
  • (M3 & M4): optimization based Gaussian graphical model estimation methods Cai et al., 2016a and Lee and Liu, (2015), respectively
Methods (M0) and (M1) are based on the matrix-variate Gaussian model, while the others are designed for the vector-variate model. Since the vector-variate model does not account for temporal correlation, methods (M2), (M3), and (M4) are at a disadvantage when applied to matrix-variate data. To ensure a fair comparison, we used the whiten function from the whitening R package to whiten the simulated data before applying the vector-variate methods.
The edge detection performance of the methods was evaluated using receiver operating characteristic (ROC) curves. An ROC curve plots the true positive rate (TPR) against the false positive rate (FPR) as the detection threshold is varied by adjusting specific hyperparameters. For Methods (M0) and (M2), which detect non-zero spatial edges based on rejecting edge-wise null hypotheses (Eq. 5), we generated ROC curves by varying the $p$-value threshold while keeping other hyperparameters fixed. For the sparsity hyperparameter of Method (M0), we fixed $\gamma_{i}=1\text{e}-4$ for all $i$. For the remaining methods, ROC curves were plotted by adjusting the sparsity-controlling hyperparameter, with all other hyperparameters held constant.
The results with $m=5$, $n=5$, and four different pairs of $(p,q)$ are shown in Fig. 2. The ROC curves indicate that our method consistently outperformed the other methods, thanks to the efficient use of spatial observations across sessions. In an additional simulation study with varying $\gamma_{i}$ values, our method demonstrated moderate sensitivity to the choice of the group lasso hyperparameter, while it consistently outperformed the baseline methods across the range of tuning parameters considered, particularly in settings with high temporal and spatial dimensions.
Figure 2: Simulation results are shown for different graph configurations and temporal dimensions with $n=5$ and $m=5$. Rows represent different graph types, while columns correspond to varying spatial and temporal dimensions. The black curves represent our method (M0), with other colors representing baseline methods.
@@ 5.1.2 Simultaneous Test
In this section, we evaluate the coverage of the proposed bootstrap confidence interval (Eq. 8) using simulated data. The simulations were conducted with $m$ sessions of $n$ i.i.d. matrix-variate data, with temporal dimension $p=50$ and spatial dimension $q=30$. We generated 3000 bootstrap samples to construct confidence intervals for $\lVert T_{E}\rVert_{\infty}$ over two edge sets: $E_{\text{off}}=\{(i,j):i\neq j\}$ and $E_{\text{zero}}=\{(i,j):\Omega^{(\mathcal{S},l)}_{ij}=0,\ \forall l=1,\dots,m\}$.
@@
Table 1: Empirical coverage of the confidence region for $m=5,p=50,q=30$.

### v2

@@ 4 Theoretical Properties
We make the following assumptions on the multiple MGGM and the observed dataset $\mathcal{D}$.
@@
For each $l\in[m]$, let $\{\lambda^{(\mathcal{T},l)}_{i}\}_{i=1,\dots,p}$ be the eigenvalues of $\Sigma^{(\mathcal{T},l)}$ while $\frac{1}{\kappa_{3}}\leq\lambda^{(\mathcal{T},l)}_{1}\leq\lambda^{(\mathcal{T},l)}_{2}\leq\dots\leq\lambda^{(\mathcal{T},l)}_{p}\leq\kappa_{3}$ for some constant $\kappa_{3}>0$; define and assume $\{\lambda^{(\mathcal{S},l)}_{i}\}_{i=1,\dots,q}$ similarly for $\Sigma^{(\mathcal{S},l)}$.
@@
Let $d:=\max_{i}\left\lvert\left\{j\in[q]\backslash\{i\}:\Omega^{(\mathcal{S},l)}_{ij}\neq 0\text{ for some }l\in[m]\right\}\right\rvert$ be the group-wise maximum node degree. We assume group sparsity of the partial-correlation graph in spatial association by $d\cdot\frac{\max\{m,\log(mn_{0}pq)\}}{(mn_{0}p)^{1/2}}\rightarrow 0$ as $n_{0}\rightarrow\infty$.
@@
We assume the temporal precision matrix $\Omega^{(\mathcal{T},l)}$ for each $l\in[m]$, has Cholesky decomposition as in Eq. 9 where $\beta^{(\mathcal{T},l)}$ satisfies $\lvert\beta^{(\mathcal{T},l)}_{st}\rvert<\kappa_{5}(t-s)^{-\alpha_{l}-1}$ for any $t$ and $s$ such that $s<t$ and some $\alpha_{l}>0$. We further assume that, for $\alpha_{0}=\min_{l=1,\dots,m}\alpha_{l}$, $\frac{\log(mn_{0}pq)}{(n_{0}q)^{1-1/(\alpha_{0}+1)}}\rightarrow 0$ as $n_{0}\rightarrow\infty$.
Assumption 1 assumes that the sample sizes across sample groups are balanced, with $n_{0}$ representing this common level. Assumption 2 ensures identifiability. Assumption 3 is a standard assumption on eigenvalues commonly used in covariance estimation and graphical models (Cai et al., 2016b). Assumption 4 assumes that the spatial (column) precision matrices are sparse, and it imposes a constraint on the spatial dimension relative to the number of samples, the temporal dimension, and the number of graphs. In the special case of a single vector-variate graphical model with $m=1,p=1$, this assumption coincides with the necessary condition for root-consistent estimation of individual edges (Ren et al., 2015). Moreover, a larger temporal dimension $p$ can further relax the required sample size. The first part of Assumption 5 is reasonable for neural time series, as these data are often modeled as a low-order autoregressive process, a common assumption in the literature (Bickel and Levina, 2008; Liu and Ren, 2020). The second part, similar to Assumption 4, restricts the temporal dimension. Likewise, increasing the spatial dimension $q$ can further relax the sample size requirement.
@@ 4.1 Non-asymptotic error bound for the group Lasso estimate
We first provide a theoretical justification for our group Lasso procedure proposed in Section 3.1. Although with correlated rows, our results below demonstrate that the optimal convergence rates for the estimation error, defined as $\Delta^{(\mathcal{S},l)}_{\cdot,i}:=\widehat{\beta}^{(\mathcal{S},l)}_{\cdot,i}-{\beta}^{(\mathcal{S},l)}_{\cdot,i}$, and prediction can be still obtained compared to the case with i.i.d. samples. The proof is provided in Section C.3.
@@ 4.2 Theoretical justification for the simultaneous edge testing
We now present the theoretical results for the simultaneous test in Section 3.2. In the proof of Proposition 4.2, we show that the error in the partial correlation estimate $\widehat{\rho}^{(\mathcal{S},l)}_{ij}$ from Eq. 4 is driven by a leading term
@@
where $\widetilde{\phi}_{ij}^{(\mathcal{S},l)}:=\widetilde{\Phi}^{(\mathcal{S},l)}_{ij}-\Phi^{(\mathcal{S},l)}_{ij}=\frac{1}{n_{l}p}\sum_{k=1}^{n_{l}}\epsilon^{(\mathcal{S},k,l)\top}_{\cdot,i}\epsilon^{(\mathcal{S},k,l)}_{\cdot,j}-\Phi^{(\mathcal{S},l)}_{ij}$. Since $\widetilde{\phi}^{(\mathcal{S},l)}_{ij}$ is an average over $t\in[p]$ and $k\in[n_{l}]$, the central limit theorem ensures that the partial correlation estimate $\widehat{\rho}^{(\mathcal{S},l)}_{ij}$ and the single-edge test statistic $\widehat{T}_{ij}$ converge in distribution to Gaussian limits for each edge $(i,j)$.
The Gaussian approximation error for the multiple-edge test statistic $\lVert\widehat{T}_{E}\rVert_{\infty}$ is governed by Berry–Esseen bounds over hyper-rectangles. The seminal work of Chernozhukov et al. (2013) established these bounds in high-dimensional settings, where the dimension can exceed the sample size. Building on this, we use the near-optimal convergence rate for the Berry–Esseen bound from Chernozhukov et al. (2023) to derive a sharp Gaussian approximation error bound for $\lVert\widehat{T}_{E}\rVert_{\infty}$. The proof is provided in Section C.4.
@@
$\sup_{x>0}\left\lvert\mathbb{P}[\lVert\widehat{T}_{E}\rVert_{\infty}>x]-\mathbb{P}[\lVert T_{E}+Z\rVert_{\infty}>x]\right\rvert$
@@
The following theorem mirrors the previous proposition, with the key difference being that we replace the population covariance $S_{EE}$ with its plug-in estimator $\widehat{S}_{EE}$. In essence, as long as $S_{EE}$ is well-estimated under the $\lVert\cdot\rVert_{\infty}$-norm, the Gaussian approximation results remain valid. Using the convergence rate of $\widehat{S}_{EE}$ in the $\lVert\cdot\rVert_{\infty}$-norm, we derive the following bootstrap theorem, based on the relationship between convergence rate and bootstrap error as given in Lemma 2.1 of Chernozhukov et al. (2023).
@@
$\sup_{x>0}\left\lvert\mathbb{P}[\lVert\widehat{T}_{E}\rVert_{\infty}>x]-\mathbb{P}[\lVert T_{E}+\widehat{Z}\rVert_{\infty}>x|\mathcal{D}]\right\rvert$
@@
The above theorem establishes the theoretical foundation for the simultaneous multiple edge testing procedure in Algorithm 1. Next, we formally state the validity of our testing procedure as well as a power analysis. The proof is given in Section C.6
@@
and $\frac{1}{q}((\log\lvert E\rvert\log(n_{0}q))^{2}\log(mn_{0}pq))^{1+1/(2\alpha_{0}+1)}$. Under the null $H_{0,E}$ and Assumptions 1 to 5, the confidence region $\mathcal{C}_{E}(1-\alpha)$ estimated based on $\gamma_{i}$’s given in Theorem 4.1 satisfies $\mathbb{P}[\mathbf{0}\notin\mathcal{C}_{E}(1-\alpha)]\overset{p}{\rightarrow}\alpha.$ On the other hand, as an alternative, if
then under the same assumptions and $\gamma_{i}$’s, we have $\mathbb{P}[\mathbf{0}\notin\mathcal{C}_{E}(1-\alpha)]\overset{p}{\rightarrow}1$.
Remark 4.5.
@@
Moreover, when the temporal covariance matrices satisfy $\Sigma^{(\mathcal{T},l)}=I$ for all $l\in[m]$, the model simplifies to the multiple vector-variate Gaussian graphical model studied in Ren et al. (2019), with an effective sample size of $n_{0}p$. According to their Theorem 2.3, the optimal separating rate for detecting a single edge, measured in terms of the sum of partial correlations, is $\sqrt{m/(n_{0}p)}$. This comparison indicates that our test achieves nearly optimal performance in terms of the testable region boundary. We further conjecture that the extra $\sqrt{\log q}$ factor is intrinsic when the edge set size $|E|$ grows at a polynomial rate in $q$.
@@ 5 Numerical Studies
In Section 5.1, we present simulation studies to validate our theoretical results of Section 4 and evaluate the performance of our inference procedure for multiple matrix-variate graphical models in comparison to several baseline approaches. In Section 5.2, we apply our method to neural recordings from two Utah arrays implanted in monkey brains during a spatial working memory task. The implementation of our proposed method is available in the mmge R package, along with vignettes for reproducing our results, accessible at github.com/HeejongBong/mmge.
@@ 5.1 Simulation Studies
We evaluated the performance of our method under three spatial graph structures, illustrated in Figure 1: (1) a random graph, where edges between nodes were generated with probability $\sqrt{\frac{3}{q}}$ for each pair $(i,j)$; (2) a hub graph, where the nodes were divided into $\lceil\frac{q}{20}\rceil$ hub groups; and (3) a chain graph. Given each spatial graph structure, we generated the spatial precision matrix by sampling the non-zero entries from $\text{Unif}(0,\frac{0.3}{2^{l-1}})$, independently across sample groups $l\in[m]$. For the temporal precision matrix, we generated $\Sigma^{(\mathcal{T},l)}$ using Eq. 9. Following Assumption 5, we set $\beta^{(\mathcal{T},l)}_{st}=\kappa_{5}(t-s)^{-\alpha_{l}-1}$ for $1\leq s<t\leq p$, with $\kappa_{5}=0.2$ and $\alpha_{l}=1$. For $\Phi^{(\mathcal{T},l)}$, we used the $p\times p$ identity matrix for all $l\in[m]$.
In Section 5.1.1, we compare our method’s performance with existing methods in terms of edge detection and precision matrix estimation. In Section 5.1.2, we assess the coverage of the proposed bootstrap confidence region for $T_{E}$.
@@ 5.1.1 Edge-wise Estimation Comparison
  • (M1): matrix-variate Gaussian multi-graph estimation method by Zhu and Li (2018)
  • (M2): regression-based Gaussian graph estimation method by Ren et al. (2019)
  • (M3 & M4): optimization based Gaussian graph estimation methods by Cai et al. (2016a) and Lee and Liu (2015), respectively
Methods (M0) and (M1) are based on the matrix-variate Gaussian graphical model, while the others are designed for the vector-variate model. Since the vector-variate model does not account for temporal correlation, methods (M2), (M3), and (M4) are at a disadvantage when applied to spatiotemporal data. To ensure a fair comparison, we used the whiten function from the whitening R package to whiten the simulated data before applying the vector-variate methods.
The edge detection performance of the methods was evaluated using receiver operating characteristic (ROC) curves. An ROC curve plots the true positive rate (TPR) against the false positive rate (FPR) as the detection threshold is varied by adjusting specific hyperparameters. For Methods (M0) and (M2), which detect non-zero spatial edges based on rejecting edge-wise null hypotheses (Eq. 5), we generated ROC curves by varying the $p$-value threshold while keeping other hyperparameters fixed. For the sparsity hyperparameter of Method (M0), we fixed $\gamma_{i}=1\text{e}-4$ for all $i$. For the other methods, ROC curves were plotted by varying the sparsity hyperparameter, with other hyperparameters held constant.
The results with $m=5$, $n=5$, and four different pairs of $(p,q)$ are shown in Fig. 2. The ROC curves indicate that our method consistently outperformed the other methods, thanks to the efficient use of spatial observations across sample groups. In an additional simulation study with varying $\gamma_{i}$ values, our method demonstrated moderate sensitivity to the choice of the group Lasso hyperparameter, while it consistently outperformed the baseline methods across the range of tuning parameters considered, particularly in settings with high temporal and spatial dimensions.
Figure 2: Simulation results are shown for different graph configurations and temporal dimensions with $n=5$ and $m=5$. Rows represent different graph types, while columns correspond to varying spatial and temporal dimensions. Curves represent our method (M0) and the other baseline methods.
@@ 5.1.2 Simultaneous Test
In this section, we evaluate the coverage of the proposed bootstrap confidence interval (Eq. 8) using simulated data. The simulations were conducted with $m$ sample groups of $n$ i.i.d. matrix-variate data, with temporal dimension $p=50$ and spatial dimension $q=30$. We generated 3000 bootstrap samples to construct confidence intervals for $\lVert T_{E}\rVert_{\infty}$ over two edge sets: $E_{\text{off}}=\{(i,j):i\neq j\}$ and $E_{\text{zero}}=\{(i,j):\Omega^{(\mathcal{S},l)}_{ij}=0,\ \forall l=1,\dots,m\}$.
@@
Table 1: Empirical coverage of the confidence region when $m=5,p=50,q=30$.